OpenStudy (anonymous):

Evaluate \[\sum_{k=0}^{\infty} \frac{2k + 1}{(4k + 1)(4k+3)(4k+5)}\]

OpenStudy (anonymous):

How are you gonna do it without, Mathematica or Wolfram Alpha?

OpenStudy (anonymous):

Use partial fractions to break the fraction up.

OpenStudy (anonymous):

I am trying partial fractions :-/

OpenStudy (anonymous):

Is that the only way?

OpenStudy (anonymous):

most likely yes. It might turn into a telescoping sum.

myininaya (myininaya):

i would like to try this problem, but my nails aren't dry yet. so its hard to type

OpenStudy (anonymous):

umm I am getting \[\sum_{k=0}^{\infty}\frac{2k+1}{16}\times\left(\frac{1}{4k+1} + \frac{2}{4k+3} - \frac{3}{4k+5} \right)\]

OpenStudy (anonymous):

im getting this:\[\frac{2k+1}{(4k+1)(4k+3)(4k+5)}=\frac{1}{16}\left(\frac{1}{4k+1}\right)+\frac{1}{8}\left(\frac{1}{4k+3}\right)-\frac{3}{16}\left(\frac{1}{4k+5}\right)\]which isnt helping too much... i cant really afford to spend too much time on this as i have my own finals to study for, but here are some small things that may or may not lead to a solution: Note that:\[\sum_{k=0}^{\infty}\frac{1}{4k+1}=1+\sum_{k=0}^{\infty}\frac{1}{4k+5}\] Also note that:\[\frac{(4k+1)+(4k+5)}{2}=4k+3\]

OpenStudy (anonymous):

Thanks Joe! and good luck for your exams!

myininaya (myininaya):

i wish it was a telescoping series :( What ways have you learned ishaan?

OpenStudy (anonymous):

i don't know, a friend of mine gave me this problem; i only know the basic stuff arithematic, geometric and harmonic, series of square and cubes.